SHOW ALL WORK!!!

?the addition of hydrochoric acid to silver nitrate solution precipitate silver chloride according to the following reaction:AgNO_3 (aq)+HCl(aq) \rightarrow AgCl(s)+HNO_3(aq)

when 50.0 ml of 0.1 M AgNO_3 is combined with 50.0 ml of 0.1 M HCl in a coffee cup calorimeter, the tempertature changes from 23.4 to 24.21*C. Calculate the \Delta H_r_x_n for the reaction per mole of HCl. Use 1.0 g/ml as the dnsity of the solution and 4.184 j/g*C as the specific heat capacity.
Answer #1

V = 50 ml

M = 0.1 AgNO3

V2 = 50 ml

M2 = 0.1 M HCl

dT = 24.21 - 23.4 = 0.81

Hrxn = Q/n

Q = m*Cp*dT

Q = (50+50)*4.184*(0.81) = 338.904 J

mol Hcl = MV = 50*0.1 = 5 mmol = 5*10^-3 mol

then

Hrxn = (338.904 )/(5*10^-3) = 67780.8 J

Hrxn = -67.781 kJ/mol HCl

its negative since the reaction is exothermic